The length of own leg of a right triangle is 12 Meters. The length of the hypotenuse is 6 meter longer then the other leg. Find the lengths of the hypotenuse and the other leg
Answer by ewatrrr(24785)

 
Hi,
The length of the hypotenuse is 6 meter longer then the other leg
Let x and (x+6)represent the length of the other leg and the hypotenuse
Applying the Pythagorean Theorem:
12^2 + x^2 = (x+6)^2
12^2 + x^2 = x^2 + 12x+ 36
12x = 108
x = 9m, the other leg and the hyotenuse is 15m
CHECKING our Answer***
144 + 81 = 225